OpenStudy (anonymous):

Solve the system for the x-coordinate of the point of intersection. 3x + y = -2 2x - 5y = 4

OpenStudy (anonymous):

So then all that you're looking for is the x-coordinate of the intercept?

OpenStudy (anonymous):

3x+y=-2 2x-5y=4 I would multiply the top equation by 5.

OpenStudy (anonymous):

3(5)x+(5)y=-2(5) 2x-5y=4

OpenStudy (anonymous):

15x+5y=-10 2x-5y=4 Now you can add the equations together.

OpenStudy (anonymous):

17x=-6 and solve for x by dividing by 17 on both sides

OpenStudy (anonymous):

It certainly isn't a nice number, but -6/17 seems to be the answer.

OpenStudy (anonymous):

i got it thanx alot i love how u explain nice work